OpenStudy (anonymous):

Help pls!! Set \(\overline{N}:=N\cup\{\infty\}\), where \(\infty\) is a "point" not in \(N\). Define\(d\): \( \overline{N}x\overline{N}\rightarrow \mathbb{R}\) by\\\\ \(d(m,n):= |\frac{1}{n} - \frac{1}{m} |\) \(d(n,\infty) = d(\infty, n):= \frac{1}{n} \) \(d(\infty,\infty) := 0 \) Show that a) \(d\) defines a metric on\(\overline{N}\)

OpenStudy (kinggeorge):

Let's see then. To be a metric, we have a few conditions we need to satisfy. First off, is \(d\) always non-negative?

OpenStudy (anonymous):

hmm i am not sure its not given, or should we understand it from question context?

OpenStudy (kinggeorge):

Well, the absolute value of anything is always non-negative, and since both 0 and \(1/n\) are non-negative, we know that \(d\) is non-negative.

OpenStudy (anonymous):

ok... very good explanation

OpenStudy (kinggeorge):

Next, we need to show \(d(m, n)=0\) iff \(m=n\). First, if \(m=n\) it's either \[|\frac{1}{m}-\frac{1}{m}|=|0|=0\]or \(d(\infty,\infty)=0\).

OpenStudy (anonymous):

ok..

OpenStudy (kinggeorge):

If \(m\neq n\), we just have to concern ourselves with \[|\frac{1}{n}-\frac{1}{m}|=|\frac{m-n}{nm}|\]However, since \(n\neq m\), this is non-zero. Hence, \(d(m, n)=0\) iff \(m=n\).

OpenStudy (kinggeorge):

Third, we need to show that \(d(m, n)=d(n ,m)\). If either \(m ,n\) are infinity, this is given to us in the definition of \(d\). However, note that \[|\frac{1}{n}-\frac{1}{m}|=|-\left(\frac{1}{m}-\frac{1}{n}\right)|=|\frac{1}{m}-\frac{1}{n}|\]So we've proven that part.

OpenStudy (anonymous):

hm is it enough for this question or do we need to show more things ?

OpenStudy (kinggeorge):

Finally, we need to show that \(d(m, o)\leq d(m, n)+d(n ,o)\) for all \(m ,n, o\in\mathbb{N}\)

OpenStudy (anonymous):

ok..

OpenStudy (kinggeorge):

Fortunately, we know that \[|x+y|\leq|x|+|y|\]for all real numbers, so this means that \[\left|\frac{1}{o}-\frac{1}{m}\right|\leq\left|\frac{1}{n}-\frac{1}{m}\right|+\left|\frac{1}{o}-\frac{1}{n}\right|\]Since \[\frac{1}{n}-\frac{1}{m}+\frac{1}{o}-\frac{1}{n}=\frac{1}{o}-\frac{1}{m}\]

OpenStudy (kinggeorge):

Now you also need to include a case about infinities, where one or two of the numbers is infinity.

OpenStudy (anonymous):

hmm..

OpenStudy (kinggeorge):

Or all three are infinity. If this is the case, it's easy to see that it's all 0, so no problem there.

OpenStudy (kinggeorge):

If \(o=\infty\), then we have \[d(m,\infty)=\frac{1}{m}\]\[d(m, n)+d(n ,o)=\left|\frac{1}{m}-\frac{1}{n}\right|+\frac{1}{n}\]which is fairly obviously greater than or equal to \(1/m\).

OpenStudy (kinggeorge):

If \(n=o=\infty\), we get that \[d(m, \infty)=\frac{1}{m}\]and \[d(m, \infty)+d(\infty, \infty)=\frac{1}{m}+0\]And \[\frac{1}{m}\leq\frac{1}{m}\]

OpenStudy (kinggeorge):

Therefore, \(d\) is a metric.

OpenStudy (anonymous):

ok cool... you have explained it very good, i will study it again next time..

OpenStudy (anonymous):

i have also part b) for same question, i will post it right now, maybe its wont take too much your time as this question...